# Numerical Methods for a 2D “Bad” Boussinesq Equation: RK4, Strang Splitting, and High-frequency Fourier Modes

### v1

@@ 6 NUMERICAL RESULTS
The numerical computation is conducted using Octave software and especially its built-in two-dimensional fast Fourier transform function fft2 and the inverse ifft2. The exact solution of (1.2) from [exact_barrera] is of the form
@@
Figure 4: The $L^{\infty}$ error for both numerical methods up to $t=40$ for the first simulation.
@@
For the first simulation, we use initial condition of the form (6.1) with $\alpha=0.6$, $\beta=0.6$, and $b_{0}=10^{6}$ over the domain $\Omega=[-70,70]^{2}$. Using this setting, we numerically solve the Boussinesq system (4.1) for $t\in[0,40]$ using both the Runge-Kutta method and Strang splitting with $\Delta x=\Delta y=0.25,\Delta t=0.1$. The exact solution with this setting is a soliton wave heading north-east direction with constant speed. Due to the automatic Fourier periodic boundary condition, waves passing through the boundary will reappear again from the opposite side. As a consequence, we only compute the $L^{\infty}$ error over the subdomain $y\in[-5,5]$: in this subdomain, the waves from periodic boundary condition will not appear at least before $t=40$. The plot of the numerical solutions are shown in Figure 2 and 3. As we can see, the effect of periodic boundary condition can be seen around the corners of the domain. However, both numerical solutions have solitons heading in the same direction as the exact solution, and with approximately the same shape as well. The error is plotted in Figure 4: we can see that both methods produce good results, with $e(40)\approx 0,005<2.8\%$ of the initial condition’s absolute peak $\max|U(x,y,0)|=0.18$. The error plot shows that the performance of Runge-Kutta 4 (RK4) is slightly better than Strang splitting for $t\in[5,40]$.
$0.99\pi^{2}\leq\frac{L_{x}^{4}k^{2}+(L_{x}L_{y})^{2}j^{2}}{L_{y}^{2}j^{4}},$
Lastly, we perform another numerical simulation on a rectangular domain for all $(x,y)\in[-40,40]\times[-20,20]$ with Dirichlet boundary condition $U(x,y,t)=0$ on the set of spatial points
@@
To implement the Dirichlet boundary condition, we define a boundary matrix B of the same size as U. Matrix B is such that all the matrix elements corresponding to spatial points $(x,y)\notin\partial\Omega_{\text{dirichlet}}$ have values 1, while the others have value 0. In each time iteration after we inverse the $\widehat{\textbf{U}}$ and $\widehat{\textbf{V}}$ to get the solution U and V, we perform element-wise multiplication $\textbf{B}\odot\textbf{U}$ and $\textbf{B}\odot\textbf{V}$, and then Fourier transform again to get the new $\widehat{\textbf{U}}$ and $\widehat{\textbf{V}}$ for Dirichlet boundary condition. The results are shown in Figure 7-9. We can see a part of the soliton has been reflected against the north wall at $t=10$ indicated by red color, and at $t=20$ longer part has been reflected. At $t=30$, we can see the approximate soliton is now heading south-east, and at $t=45$ we can see a part of it has been reflected off the south wall indicated by blue color.

### v2

@@ 6 NUMERICAL RESULTS
The numerical computation is conducted using Octave software and especially its built-in two-dimensional fast Fourier transform function fft2 and the inverse ifft2. The exact solution of (1.2) from [3] is of the form
@@
For the first setting of initial condition and domain, we use initial condition of the form (6.1) with $\alpha=0.6$, $\beta=0.6$, and $b_{0}=10^{6}$ over the domain $\Omega=[-70,70]^{2}$. The exact solution with this setting is a soliton wave heading north-east direction with constant speed. Using this setting, we numerically solve the Boussinesq system (4.1) for $t\in[0,40]$ using both the Runge-Kutta method and Strang splitting with $\Delta x=\Delta y=0.25$ and $\Delta t=0.1,0.2$. Due to the automatic Fourier periodic boundary condition, waves passing through the boundary will reappear again from the opposite side. As a consequence, we only compute the $L^{\infty}$ error over the subdomain $y\in[-5,5]$: in this subdomain, the waves from periodic boundary condition will not appear at least before $t=40$. The plot of the numerical solutions using $\Delta x=\Delta y=0.25$ and $\Delta t=0.1$ are shown in Figure 3 and 4. As we can see, the effect of periodic boundary condition can be seen around the corners of the domain. However, both numerical solutions have solitons heading in the same direction as the exact solution, and with approximately the same shape as well. The $L^{\infty}$ error is plotted in Figure 5 for both time step scenarios $\Delta t=0.1,0.2$. We can see
Figure 6: The $L^{\infty}$ error for both numerical methods up to $t=40$ for the first simulation, using $\triangle x=0.5$ and three different time steps $\Delta t=0.1,0.25,0.45$.
that both methods produce good results, with $e(40)\approx 0,005<2.8\%$ of the initial condition’s absolute peak $\max|U(x,y,0)|=0.18$. The error plot shows that the performance of Runge-Kutta 4 (RK4) is slightly better than Strang splitting for $t\in[5,40]$ for $\Delta t=0.2$. However, for $\Delta t=0.1$, the performance of both methods are almost indistinguishable. Additionally, we also simulate with $\Delta x=\Delta y=0.5$ and $\Delta t=0.1,0.25,0.45$ to compare the numerical errors. The $L^{\infty}$ errors resulting from cases $\Delta x=\Delta y=0.5$ with $\Delta t=0.1,0.25,0.45$ are shown in Figure 6. From these plots, it is clear that the performance difference between Runge-Kutta 4 and Strang splitting becomes more noticeable as $\Delta t$ increases. For $\Delta t=0.1,0.25,0.45$, the Runge-Kutta 4 performs well and shows similar results for all three time steps, while the Strang splitting becomes approximately as good as Runge-Kutta only when $\Delta=0.1$.
$0\leq L_{x}^{4}k^{2}+(L_{x}L_{y})^{2}j^{2}-0.99\pi^{2}L_{y}^{2}j^{4},$
@@
Lastly, we perform another numerical simulation on a rectangular domain $40,40]\times[-20,20]$ with Dirichlet boundary condition $U(x,y,t)=0$ on the set of spatial points
@@
To implement the Dirichlet boundary condition, we define a boundary matrix B of the same size as U. Matrix B is such that all the matrix elements corresponding to spatial points $(x,y)\notin\partial\Omega_{\text{dirichlet}}$ have values 1, while the others have value 0. In each time iteration after we inverse the $\widehat{\textbf{U}}$ and $\widehat{\textbf{V}}$ to get the solution U and V, we perform element-wise multiplication $\textbf{B}\odot\textbf{U}$ and $\textbf{B}\odot\textbf{V}$, and then Fourier transform again to get the new $\widehat{\textbf{U}}$ and $\widehat{\textbf{V}}$ for Dirichlet boundary condition. The results are shown in Figure 9-11. We can see a part of the soliton has been reflected against the north wall at $t=10$ indicated by red color, and at $t=20$ longer part has been reflected. At $t=30$, we can see the approximate soliton is now heading south-east, and at $t=45$ we can see a part of it has been reflected off the south wall indicated by blue color.